1. A product has a price elasticity (of demand) equal to 1.50. If price increases by six percent, what will be the decrease in quantity demanded?

2. A product has an income elasticity of 0.75. If income rises by 8 percent, what will be the increase in demand?

3. In question 2, is the product most likely a luxury or necessity? Why?

4. The cross price elasticity between two products, L and M, is 0.40 (that is, the change in demand for L with respect to the change in the price of M). If the price of M rises by 10 percent, by how much will the demand for L change?

5. In question 4, are L and M substitutes or complements, and why?
Answer #1

1) 1.50=D/6

D=6*1.5

D= 9%

2) 0.75=D/8

D=8*0.75

D=6%

3) The product is a necessity because the price elasticity is inelastic

4) 0.40=D/10

D=10*0.40

D=4%

5) The product is a substitute because the price elasticity is positive